government expenditure increased from 537,154,659,558 on 2019 to 559,740,919,125 on 2020, Assume Marginal Propensity to Consume is 0.80 (some studies show that MPC is between 0.7 to 0.95). Calculate:

a- Multiplier effect.

b- Increase in GDP due to increase on government expenditure from 2019 to 2020.

c- How does your answer on part b will be if there is crowding out effect? (you only need to say: increase, decrease, or it does not change)
Answer #1

1) Multiplier effect refers to the increase in income or real GDP by more than percentage rise in Government expenditure.

Multiplier = Change in income/ Change in goverment expenditure

= 1/(1-MPC)

= 1/(1-0.8)

= 1/0.2

=5


2) Increase in real GDP = Multiplier × Increase in government expenditure

= 5 × 22586259567

= 112931297835 SAR


3) If there is a crowding-out effect, the rise in income will be affected negatively due to the interest rate effect. In other words, the rise in government expenditure will cause a rise in market interest rate which will lead to a fall in private induced investment. So, a part of rising in aggregate demand (due to govt expenditure) will be compensated by a fall in investment expenditure. So, overall the real GDP or Income will decrease from the answer which has been arrived at in part b.


answered by: Dr.E